OpenStudy (anonymous):

Simplify using the Order of Operations: -3(-5) - 3 (-4)^3

OpenStudy (anonymous):

Ok, so what's the first move?

OpenStudy (anonymous):

i got -177

OpenStudy (anonymous):

( )

OpenStudy (anonymous):

oh, you just want to check answer.. No, it's definitely not negative

OpenStudy (anonymous):

so is it just 177

OpenStudy (anonymous):

no

OpenStudy (mathmagician):

=-3(-5-(-4)^3)=3(5+4^3)=32*69=177

OpenStudy (anonymous):

Am I missing something? \[-3(5) - 3(-4)^3\] Is the expression we're working with right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, so which part are you going to start with?

OpenStudy (mathmagician):

sorry, the end is 3*69=207

OpenStudy (anonymous):

after i redid it that is what I got thanks again